# Deep Learning based Key Information Extraction from Business Documents: Systematic Literature Review

### v1

@@ 5.1. Overview
The evolution of model sizes in terms of their number of parameters (in millions) is shown in figure 5. Note the necessary logarithmic scale of the y-axis due to the approach by (He et al., 2023b), which is based on GPT3 and therefore includes 175 billion parameters. The orange line indicates the trend over time, which shows a relatively small increase. Based on the analyzed approaches, the median number of trainable parameters is 157 million. Most of the approaches tend to have between 20 and 500 million parameters, which did not change significantly between 2017 and 2023. This is in stark contrast to the research field of LLMs in general, where models continue to grow exponentially with sizes often in the high billions.
@@ 5.2.1. Sequence-based methods
Table LABEL:sequencetab of the electronic supplementary material presents the results regarding the analysis of sequence-based approaches. We also include hybrid KIE approaches that are based on multiple paradigms (e.g., a combination of graph-based and sequence-based methods) in this analysis and present the findings regarding the sequence-based subsystem. Most of the methods are defined on word-level (n=40), i.e., each word of a document image represents one element in the sequence. In some cases, a segment or sentence-level granularity is chosen (Gu et al., 2021; Li et al., 2021b, c; Ning et al., 2021; Tang et al., 2021; Wang et al., 2021a, 2022a; Kuang et al., 2023; Tu et al., 2023; Zhang et al., 2023a), which considers sentences as semantic entities and subsequently calculate aggregated visual and textual embeddings. Other possible granularities are on character-level (Guo et al., 2019; Yu et al., 2020), token-level (Wang et al., 2021a) or cell-level (Li et al., 2021a). Also, (Oral et al., 2020; Wang et al., 2021a, 2022a) use two different granularity types simultaneously, which allows the model to consider both granular and coarse features.
@@
26 out of the 57 sequence-based approaches do not consider a pre-training of the models and subsequent fine-tuning steps for the KIE task. This is somewhat surprising, as it has been shown that obtaining a general document understanding through pre-training is a promising research direction. However, it should be noted that pre-training has been increasingly studied in the analyzed work since 2022. If pre-training is conducted, the average number of documents used is around 6 million. This highlights the relatively large amount of data that is required to pre-train corresponding models. The largest document corpus for pre-training is used by (Cao et al., 2022a), which consists of 43 million documents. Deviating from this is the approach of (Nguyen et al., 2021), which only uses 5,170 documents for pre-training procedures, while still achieving competitive benchmark results. Pre-training procedures are usually very resource-intensive. For example, (Xu et al., 2020) report that their largest model variant required 170 hours to finish one training epoch of a dataset consisting of 6 million documents. In around half of the cases, the IIT-CDIP dataset is used. This dataset has been proposed in (Lewis et al., 2006) and includes documents of the Legacy Tobacco Document Library [^4] https://www.industrydocuments.ucsf.edu/tobacco/ regarding lawsuits against the tobacco industry in the 1990s. Another common dataset is RVL-CDIP, which however represents a subset of IIT-CDIP and therefore includes no additional documents. Also, in six cases a private document collection has been used for pre-training procedures (Hwang et al., 2021a; Nguyen et al., 2021; Arroyo et al., 2022; Kim et al., 2022; Lee et al., 2022; Davis et al., 2023). Besides these common datasets, there exist other large document collections such as DocBank, which is however only considered by (Li et al., 2021c, 2022c; Tang et al., 2023; Yang et al., 2023a).
Table LABEL:sequencetab also lists the employed pre-training tasks, with the most common being Masked Visual-Language Modeling (MVLM) proposed by (Xu et al., 2020). This task is strongly related to the more general Masked Language Modeling (MLM), which was originally introduced with BERT (Devlin et al., 2019). For the sake of MVLM, random tokens of the input sequences are masked where the goal of the model is to predicted the masked tokens given the surrounding context. Importantly, positional information (such as word coordinates in the document image) is not masked and thus helps the model with the reconstruction. In this way, the model learns language contexts while also exploiting positional information. Various other related tasks have been proposed over time, which also explicitly include image modalities. This allows KIE models to fully exploit textual, positional and visual information during pre-training. When authors propose end-to-end and/or generative KIE methods, specific pre-training objectives are defined in order to learn text reading and producing outputs in an autoregressive manner. See also chapter 6.1 for a further discussion.
Regarding the distribution of integrated encoder and decoder architectures, 11 of the 57 sequence-based methods do not incorporate a dedicated textual encoder, while 17 of the methods do not use a visual encoder. In (Palm et al., 2017; He et al., 2023b), neither types of encoders are employed. In most cases, BERT or derived variants such as RoBERTA (Liu et al., 2019b) and SBERT (Reimers and Gurevych, 2019) are used for textual encoding. This is not surprising, given the popularity of BERT-based models in NLP. Another common encoding method is the LayoutLM-family, more specifically LayoutXLM, LayoutLM and LayoutLMv2, used by (Appalaraju et al., 2021; Nguyen et al., 2021; Zhang et al., 2021; Cao et al., 2022a; Du et al., 2022; Gu et al., 2022; Li et al., 2022c; Wang et al., 2022a; Deng et al., 2023a; He et al., 2023c; Li et al., 2023b). This can be explained by the popularity of these models in the context of DU as well as their extensive pre-training, which allows approaches to utilize expressive embeddings. Some authors also make use of more traditional embedding methods such as N-gram-models (Gbada et al., 2023), Word2Vec (Zhang et al., 2021) or FastText (Oral et al., 2020).
Regarding visual encoders, one can see that there is no clear dominance of one particular model. Nonetheless, the most common visual backbone category is represented by ResNet models (He et al., 2016), which are also popular in related CV tasks. ResNet50 and ResNet18 are the most often used variants of this architecture (Yu et al., 2020; Appalaraju et al., 2021; Gu et al., 2021; Li et al., 2021c; Wang et al., 2021b; Cao et al., 2022b; Liao et al., 2023; Yu et al., 2023), however other variants that have been proposed as improvements such as ResNeXt101-FPN and ConvNeXt-FPN are employed as well (Xu et al., 2021; Gu et al., 2022; Dhouib et al., 2023; Luo et al., 2023; Yang et al., 2023a; Yu et al., 2023). Also, (Kim et al., 2022; Davis et al., 2023; Dhouib et al., 2023) incorporate the more recent Swin Transformer (Liu et al., 2021) as the visual encoder, which captures local and global image features through sequential processing of non-overlapping image patches and shifted window self-attention mechanisms. Interestingly, only the approach by (Guo et al., 2024) integrates a Vision Transformer (ViT). This model also leverages transformers for visual tasks, however compared to Swin Transformers, the input images are processed as sequences of patches without hierarchical divisions. Document Image Transformer (DiT) (Li et al., 2022b), which in turn is a self-supervised improvement over ViT, is also only used in the approach by (Huang et al., 2022).
@@
The distribution of chosen decoders shows that the majority of approaches use sequence labeling layers for decoding. This usually consist of a linear and a Softmax layer, which assigns a probability distribution over all possible fields to each token in the sequence. The task of KIE is then performed by choosing the field with the highest probability according to the probability distribution. Besides, some approaches alternatively use LSTMs (Hochreiter and Schmidhuber, 1997), BiLSTMs (Schuster and Paliwal, 1997), Conditional Random Fields (CRF) (Lafferty et al., 2001) or a combination thereof (Palm et al., 2017; Guo et al., 2019; Oral et al., 2020; Sage et al., 2020; Yu et al., 2020; Hamdi et al., 2021; Klaiman and Lehne, 2021; Ning et al., 2021; Wang et al., 2021a, b; Kuang et al., 2023; Zhang et al., 2023a). BiLSTMs are often used as they effectively capture contextual information in a bidirectional fashion. CRFs are useful, as they model the dependencies between labels in a sequence and incorporate a global context. These more complex models can therefore provide additional information compared to a simple Softmax layer and can potentially increase the extraction performance. When looking at the decoder choice over time, one can see that the two aforementioned variants have not been used as frequently since 2022. Also, especially in 2023, newer decoding methods were examined. Some authors also use existing KIE approaches for decoding. For example, (Hong et al., 2022) employ the approach proposed by (Hwang et al., 2021b) as the decoding mechanism and (Deng et al., 2023a) base their work on the system by (Kim et al., 2022). Besides the already mentioned architectures, a plethora of different models are employed. We refer to table LABEL:sequencetab for a listing of the methods.
@@
All in all, it can be stated that there are many different approaches with regard to the employed encoder and decoder architectures. BERT-variants are most often used for textual encoding, while ResNet-based models are used as visual backbones. Decoding steps are usually performed either by simple sequence labeling layers or more complex architectures that allow to incorporate certain dependencies within the token sequences.
@@ 5.2.2. Graph-based methods
The findings regarding the analyzed graph-based approaches are presented in table LABEL:graphtab of the electronic supplementary material. A key property in which the approaches differ, is how the underlying graph representation is constructed. 10 of the approaches use a fully connect graph (Liu et al., 2019a; Carbonell et al., 2020; Luo et al., 2020; Tang et al., 2021; Zhang et al., 2021; Wang et al., 2022a; Zhang et al., 2022a, b; Gemelli et al., 2023; Zhang et al., 2023b), which means that every single node of the graph is connected to each other. As a result, the resulting graphs usually have a very large number of edges. In 8 cases, a k-nearest neighbors algorithm is chosen, where each node only has k neighboring nodes. The approaches use different values for k, namely 4 (Zhang et al., 2022b; Belhadj et al., 2023a; Hamri et al., 2023), 5 (De Trogoff et al., 2022), 8 (Belhadj et al., 2023a), 10 (Carbonell et al., 2020), 15 (Deng et al., 2023b) or even 36 (Zhang et al., 2023c). Other common methods are to identify the nearest neighbors in the four major directions up, down, left and right (Lohani et al., 2019; Qian et al., 2019; Gal et al., 2020; Krieger et al., 2021; Li et al., 2023b) or using the ß-skeleton graph algorithm, which connects nodes based on their geometric proximity depending on the parameter ß, where all approaches choose ß=1 (Lee et al., 2021, 2022, 2023). (Holeček, 2021; Wang et al., 2023a) divide documents into distinct areas starting from each node (e.g., in 45 degree angles) and identify closest neighboring nodes in each of these segments. (Cheng et al., 2020) emit 36 rays out of each node and declare all other nodes, whose bounding box is crossed by these rays, as a neighbor. In (Yu et al., 2020; Davis et al., 2021; Hwang et al., 2021b; Gemelli et al., 2023; Wang et al., 2023a), the graph creation is iteratively learned by a neural network and thus not defined by a deterministic algorithm.
Regarding different methods, how the information is propagated through the graphs, in most cases (n=10), a regular Graph Convolutional Network (GCN) is employed (Lohani et al., 2019; Qian et al., 2019; Gal et al., 2020; Luo et al., 2020; Wei et al., 2020; Davis et al., 2021; Holeček, 2021; Lee et al., 2022; Wang et al., 2022a; Lee et al., 2023). GCNs propagate information between the graph nodes to iteratively learn representations, which capture both local and global contexts. Each iteration (i.e., graph convolution) calculates new embeddings for the nodes and edges. In (Liu et al., 2019a; Yu et al., 2020; Tang et al., 2021; Zhang et al., 2021; Deng et al., 2023b), a variation of a GCN, which is defined on node-edge-node-triplets, is used in conjunction with a Multilayer perceptron (MLP). In this setting, the features of the node itself, the edge and all neighboring nodes are fused in order to obtain the new node embedding. Another common method for graph propagation is using Graph Attention Networks (GATs), either with our without multi-head attention (Carbonell et al., 2020; Hua et al., 2020; Belhadj et al., 2021; Krieger et al., 2021; De Trogoff et al., 2022; Zhang et al., 2022a, b; Belhadj et al., 2023a, b; Zhang et al., 2023c). This architecture utilizes the attention mechanism based on transformers to calculate different weight coefficients for each neighboring node, allowing the model to focus on the most relevant information during graph propagation and thus improving the overall node representations (Veličković et al., 2018). For further details regarding propagation types, we refer to the listing in table LABEL:graphtab.
@@
With respect to the graph propagation, one can also analyze the number of propagation layers. The number of layers not only determines the overall complexity of the model, but also the receptive field of each node, determining how far it can aggregate information from its neighboring nodes. Each layer increases the depth of this receptive field by one. On average, the approaches consist of 3.6 layers. (Liu et al., 2019a) conducted ablation studies regarding different numbers of layers and found that the optimal number of GCN layers is 2, which however can be task-dependent and should be considered on a case-by-case basis. With regard to the GAT-based approaches, one need to decide on the number of attention heads. (Belhadj et al., 2021) conducted a study with different attention heads and came to the conclusion that in their case, 26 attention heads produced the best results. The authors also mention that correlations between the number of attention heads and the number of fields to be extracted might exist. Here, too, it is necessary to examine which variant works best in each individual case.
@@
The approaches use different methods how KIE is performed, i.e., how the graph structure is used to extract information from documents. Most of the approaches use a simple node classification (Lohani et al., 2019; Gal et al., 2020; Belhadj et al., 2021; Davis et al., 2021; Krieger et al., 2021; Lee et al., 2021; Belhadj et al., 2023a, b; Hamri et al., 2023), i.e., after the embeddings of each graph node have been obtained, they are used to classify each node into one of the fields to be extracted. Also common are BiLSTMs, CRFs, their combination in the form of BiLSTM-CRFs (Liu et al., 2019a; Qian et al., 2019; Hua et al., 2020; Yu et al., 2020; Tang et al., 2021; De Trogoff et al., 2022; Zhang et al., 2022b; Deng et al., 2023b) as well as sequence labeling layers (Wei et al., 2020; Wang et al., 2022a; Li et al., 2023b; Zhang et al., 2023c). Besides, multiple other methods are used to perform the KIE task. For example, (Lee et al., 2022) use the Viterbi algorithm, which identifies the most likely sequence of states by iteratively calculating probabilities and backtracking through the graph.
@@
In by far the most cases (n=18), the graphs are defined on word-level, which means that every word on a document image represent one node in the constructed graph. This is also inline with the granularities for the sequence-based methods. Another common way to define graphs is to consider entire segments or sentences (n=13). The resulting graphs are therefore rather coarsely defined. 4 of the approaches consider entire text lines as nodes, regardless of whether the contained text elements are related to each other or not (Davis et al., 2021; Tang et al., 2021; Zhang et al., 2022b; Li et al., 2023b). (Wang et al., 2022a) and (Zhang et al., 2022b) define the graphs on two different granularities simultaneously, with one being more fine-granular (either words or even individual token) and the other being more coarse-granular (text segments or text lines).
@@
Another central component of graph-based methods is the definition of node features. These features determine, which types of information the approach can utilize for the KIE task. The analysis has shown that various different features have been proposed over time. In general, the node features usually consist of position-oriented values such as normalized coordinates of the node within the document image. In (Lohani et al., 2019; Luo et al., 2020; Hwang et al., 2021b), relative distances to neighboring nodes are also part of the node features. Another common feature type is related to the textual content represented by the node, e.g., word embeddings. The approaches differ in terms of which embeddings models are used. In this regard, BERT-based-embeddings are often employed, which was also the observation regarding textual encoders of sequence-based approaches. Some other methods are Byte Pair Encoding (Sennrich et al., 2016), BiLSTM and LayoutLM. In some cases, multiple text embeddings are used simultaneously, for example (Zhang et al., 2021) employ Word2Vec, BERT and LayoutLM for the word embeddings. On top of positional and textual embeddings, many approaches also integrate visual features, often obtained from models such as ResNet or LayoutLM and their variants. Besides the previously mentioned features, many authors also use hand-crafted miscellaneous features. The goal of these features is usually to incorporate certain node characteristics as well as relationships between nodes that should aid the model in performing KIE. (Lohani et al., 2019) for example include various boolean features whether a graph node represents a date, zipcode or known city, among others.
The approaches also differ regarding the size of the node feature vectors. Although in many cases the actual number is not reported in the manuscripts, on average, the features vectors have a length of around 550. This is somewhat smaller compared to common embeddings sizes such as 768, which are used by the original BERT model (Devlin et al., 2019). In general, there is little discussion about embeddings sizes for node features in related work, from which it can be concluded that these have no significant influence on the performance of the KIE systems. Much more important seem to be the different types of features which are being integrated. Another distinguishing factor of graph-based approaches is, whether they incorporate a global node in the defined graph representation. This global node can contain information about the entire document image and can be connected to the individual nodes of the graph. To this end, only the approaches by (Hua et al., 2020; Zhang et al., 2023c) use such global nodes. (Li et al., 2023b) also construct a global node, however it is not used as a information carrier and is rather formally required due to the chosen concept based on document layout trees.
@@
Edges represent the second integral part of the graph structures. It is noticeable that a large portion of work (n=17) does not incorporate any edge features. Consequently, less attention is paid to this type of feature compared to node embeddings. In the other cases, edge features are usually derived from pairwise positional relationships between the connected nodes. Most often, horizontal and/or vertical distances, both in absolute and relative values, are considered (Liu et al., 2019a; Wei et al., 2020; Tang et al., 2021; Zhang et al., 2021). The authors highlight the importance of the visual distances between two nodes (Liu et al., 2019a). The aspect ratio of the bounding boxes that span the graph nodes is also often included as edge features (Liu et al., 2019a; Yu et al., 2020; Tang et al., 2021; Lee et al., 2022, 2023). In some papers, certain edge types are defined as edge features. For example, (Qian et al., 2019) use the the edge direction (left-to-right, right-to-left, up-to-down, down-to-up) as one feature value. In only three of the graph-based approaches are visual embeddings part of the edge features (Davis et al., 2021; Lee et al., 2021, 2023). Regarding the distribution of edge directions, i.e., whether the edges are directed or undirected, there is no clear dominant variant, however in the majority of the cases (n=19), the edges are undirected. (Tang et al., 2021; Zhang et al., 2021, 2022a) did not explicitly state which type they use and it is also not clear from the manuscript. In general, the choice of the edge definition does not seem to have a significant influence on the performance of the corresponding KIE systems. Additional research is required to appropriately estimate the impact of certain design features such as the edge direction on the extraction performance.

### v2

@@ 5.1. Overview
The evolution of model sizes in terms of their number of parameters (in millions) is shown in figure 4. Note the necessary logarithmic scale of the y-axis due to the approach by (He et al., 2023c), which is based on GPT3 and therefore includes 175 billion parameters. The orange line indicates the trend over time. Based on the analyzed approaches, the median number of trainable parameters is 170 million. Most of the approaches tend to have between 20 and 500 million parameters, which did not change significantly over time. However, with the advent of LLMs, more and more approaches with several billion parameters have been proposed, as shown in the figure.
@@ 5.2.1. Sequence-based methods
Table ST4 of the electronic supplementary material presents the results regarding the analysis of sequence-based approaches. We also include hybrid KIE approaches that are based on multiple paradigms (e.g., a combination of graph-based and sequence-based methods) in this analysis and present the findings regarding the sequence-based subsystem. Most of the methods are defined on word-level, i.e., each word of a document image represents one element in the sequence. In some cases, a segment or sentence-level granularity is chosen, which considers sentences as semantic entities and subsequently calculates aggregated embeddings. Other possible granularities – albeit rarely used – are character-level, token-level or cell-level. The word-level definition is popular because it balances semantic richness with computational efficiency. It also aligns well with the OCR outputs and pretrained transformers like LayoutLM. While more fine-grained definitions can improve structural understanding, they are less common due to increased complexity and alignment challenges. Some authors use multiple granularity types simultaneously, allowing the system to consider both granular and coarse features, which can complement each other.
@@
Around half of the sequence-based approaches do not consider a pre-training of the models and subsequent fine-tuning steps for the KIE task. This is somewhat surprising, as it has been shown that obtaining a general document understanding through pre-training is a promising research direction. If pre-training is conducted, the average number of documents used is around five million. This highlights the relatively large amount of data that is required to pre-train corresponding models. The largest document corpus for pre-training is used by (Cao et al., 2022a), which consists of 43 million documents. Deviating from this is the approach of (Nguyen et al., 2021), which only uses 5,170 documents for pre-training procedures, while still achieving competitive benchmark results. Pre-training is usually very resource-intensive. For example, (Xu et al., 2020) report that their largest model variant required 170 hours to finish one training epoch of a dataset consisting of six million documents. In around half of the cases, the IIT-CDIP dataset is used. This dataset has been proposed in (Lewis et al., 2006) and includes documents of the Legacy Tobacco Document Library [^3] https://www.industrydocuments.ucsf.edu/tobacco/ regarding lawsuits against the tobacco industry in the 1990s. Another common dataset is RVL-CDIP, which however represents a subset of IIT-CDIP and therefore includes no additional documents. Also, in seven cases, a private document collection has been used for pre-training procedures. Besides these common datasets, there exist other large datasets such as DocBank, which are however rarely used. Table ST4 also lists the employed pre-training tasks, with the most common being Masked Visual-Language Modeling (MVLM) proposed by (Xu et al., 2020). This task is strongly related to the more general Masked Language Modeling, which was originally introduced with BERT (Devlin et al., 2019). For the sake of MVLM, random tokens of the input sequences are masked where the goal of the model is to predicted the masked tokens given the surrounding context. Importantly, positional information is not masked and thus helps the model with the reconstruction. In this way, the model learns language contexts while also exploiting positional information. Some authors transfer this this idea to the image modality and employ Masked Image Modeling, where the model needs to reconstruct masked image tokens. However, a limitation of this task might be the fact that document images for the large part consist of white background and thus the model tends to simply predict white pixels without considering surrounding text and image tokens (Siyuan et al., 2024). Various other related tasks have been proposed over time, which also explicitly include image modalities. This allows KIE models to fully exploit textual, positional and visual information during pre-training. When authors propose end-to-end and/or generative KIE methods, specific pre-training objectives are defined in order to learn text reading and producing outputs in an autoregressive manner. See also chapter 6.2 for a further discussion.
@@
Regarding the distribution of integrated encoder and decoder architectures, 12 approaches do not incorporate a dedicated textual encoder at all, while 25 do not use a visual encoder. In (Palm et al., 2017; He et al., 2023c), neither types of encoders are employed. In most cases, BERT or derived variants such as RoBERTA (Liu et al., 2019b) and SBERT (Reimers and Gurevych, 2019) are used for textual encoding. This is not surprising, given the popularity of BERT-based models in NLP. Another common encoding method is the LayoutLM-family, more specifically LayoutXLM, LayoutLM and LayoutLMv2. This can be explained by the popularity of these models in the context of DU as well as their extensive pre-training, which allows approaches to utilize expressive embeddings. This can also explain why many sequence-based approaches do not employ a dedicated pre-training stage, as they can make use of already pre-trained models. Some authors integrate more traditional embedding methods such as N-gram-models, Word2Vec or FastText instead. Regarding visual encoders, one can see that there is no clear dominance of one particular model. Nonetheless, the most common visual backbone is represented by ResNet models (He et al., 2016), which are also popular in related CV tasks. ResNet50 and ResNet18 are the most often used variants of this architecture, however other variants that have been proposed as improvements such as ResNeXt101-FPN and ConvNeXt-FPN are employed as well. Also, some authors incorporate the more recent Swin Transformer (Liu et al., 2021) as the visual encoder, which captures local and global image features through sequential processing of non-overlapping image patches and shifted window self-attention mechanisms. Interestingly, only the approach by (Hamri et al., 2023) integrates a Vision Transformer (ViT). This model also leverages transformers for visual tasks, however compared to Swin Transformers, the input images are processed as sequences of patches without hierarchical divisions. Document Image Transformer (DiT) (Li et al., 2022b), which in turn is a self-supervised improvement over ViT, is also only used in the approach by (Huang et al., 2022). Transformer-based visual encoders might be less commonly used due to their higher computational and training complexity, despite their potential to capture richer visual features than CNN-based models. The distribution of chosen decoders shows that the majority of approaches use sequence labeling layers for decoding. This usually consist of a linear layer and a Softmax layer, which assigns a probability distribution over all possible fields to each token in the sequence. The task of KIE is then performed by choosing the field with the highest probability according to the probability distribution. Besides, some approaches alternatively use LSTMs (Hochreiter and Schmidhuber, 1997), BiLSTMs (Schuster and Paliwal, 1997), Conditional Random Fields (CRF) (Lafferty et al., 2001) or a combination thereof. BiLSTMs are often used as they effectively capture contextual information in a bidirectional fashion. CRFs are useful, as they model the dependencies between labels in a sequence and incorporate a global context. These more complex models can therefore provide additional information compared to a simple Softmax layer and can potentially increase the extraction performance. When looking at the decoder choice over time, one can see that the two aforementioned variants have not been used as frequently since 2022. Also, especially since 2023, newer decoding methods were examined. Nonetheless, sequence labeling layers remain popular over time. Some authors also use existing KIE approaches for decoding. For example, (Deng et al., 2023a) use the model by (Kim et al., 2022) for their decoding step.
@@
Overall, it can be said that there are many different approaches to encoder and decoder mechanisms, with no dominant approach. BERT-models and layout-aware variants like LayoutLM are most commonly used for textual encoding, while ResNet-based models are used as visual backbones. Decoding steps are usually performed either by simple sequence labeling layers or by more complex architectures that allow to incorporate certain dependencies within the token sequences.
@@ 5.2.2. Graph-based methods
The findings regarding the analyzed graph-based approaches are presented in table ST5 of the electronic supplementary material. A key property in which the approaches differ, is how the underlying graph representation is constructed. Most approaches use a fully connect graph, which means that every single node of the graph is connected to each other. As a result, the resulting graphs usually have a very large number of edges. Also common is the k-nearest neighbors algorithm, where each node has k neighboring nodes. The approaches use different values for k, ranging from 4 up to 100. Choosing the value for k is no trivial task and might introduce several biases (Biescas et al., 2024). Also, increasing k will necessarily lead to more complex graph structures that include too many edges (Xu et al., 2024). As a result, KIE approaches may not be able to recognize the actual relationships within a document. Nonetheless, there is no study that investigates these aspects in depth. Other common methods are to considering neighbors in the four major directions up, down, left and right or using the $\beta$-skeleton graph algorithm, which connects nodes based on their geometric proximity depending on the parameter $\beta$. To this end, all approaches choose $\beta$=1. A few authors divide documents into distinct areas starting from each node (e.g., in 45 degree angles) and identify closest neighboring nodes in each of these segments. (Cheng et al., 2020) emit 36 rays out of each node and declare all other nodes, whose bounding box is crossed by these rays, as a neighbor. In a few KIE systems, the graph creation is not defined by a heuristic algorithm, but rather iteratively learned by a neural network. The reasoning behind this choice is that the aforementioned heuristic algorithms are not always able to adequately represent VRD layouts.
@@
Regarding different methods, how the information is propagated through the graphs, in most cases, a Graph Convolutional Network (GCN) is employed. GCNs propagate information between the graph nodes to iteratively learn representations, which capture both local and global contexts. Each iteration (i.e., graph convolution) calculates new embeddings for the nodes and edges. In some papers, a variation of a GCN defined on node-edge-node-triplets is used in conjunction with a Multilayer perceptron (MLP). In this setting, the features of the node itself, its edges and all neighboring nodes are fused in order to obtain the new node embedding. Another common method for graph propagation is using Graph Attention Networks (GATs), either with our without multi-head attention. This architecture utilizes the attention mechanism based on transformers to calculate different weight coefficients for each neighboring node, allowing the model to focus on the most relevant information during graph propagation, and thus improving the overall node representations (Veličković et al., 2018). Yet, they are less used than GCNs, which might be due to the increased complexity and computational overhead of attention mechanisms, especially when dealing with large graphs. With respect to the graph propagation, one can also analyze the number of propagation layers. The number of layers not only determines the overall system complexity, but also the receptive field of each node – determining how far it can aggregate information from its neighboring nodes. Each layer increases the depth of this receptive field by one. On average, the approaches consist of 3.7 layers. (Liu et al., 2019a) conducted ablation studies regarding different numbers of layers and found that the optimal number of GCN layers is 2, which however can be task-dependent and should be considered on a case-by-case basis. With regard to the GAT-based approaches, one needs to decide on the number of attention heads. (Belhadj et al., 2021) conducted a study with different attention heads and came to the conclusion that in their case, 26 attention heads produced the best results. The authors also mention that correlations between the number of attention heads and the number of fields to be extracted might exist. Here, too, it is necessary to examine which variant works best in each individual case.
@@
The approaches differ in the way KIE is ultimately performed. Most of the approaches use a simple node classification, i.e., after the embeddings of each graph node have been obtained, they are used to classify each node into one of the fields to be extracted using a classification layer. Also common are BiLSTMs, CRFs and their combination in the form of BiLSTM-CRFs. Besides, multiple other methods are used to perform the KIE task. For example, (Lee et al., 2022) use the Viterbi algorithm, which identifies the most likely sequence of states by iteratively calculating probabilities and backtracking through the graph.
@@
In by far the most cases, the graphs are defined on word-level, which means that every word on a document image represent one node in the constructed graph (see figure 2(a)). This is also inline with the granularities for the sequence-based methods. Another common way to define graphs is to consider entire segments or sentences. The resulting graphs are therefore rather coarsely defined. Some of the approaches consider entire text lines as nodes, regardless of whether the contained text elements are related to each other or not. A few authors define the graphs on more than one granularity simultaneously, both using fine-granular (e.g., words) and more coarse-granular (e.g., text segments) nodes.
@@
Another central component of graph-based methods is the definition of node features. These features determine, which types of information can be utilized for KIE. In general, the node features usually consist of position-oriented values such as normalized coordinates of the node within the document image. Sometimes, relative distances to neighboring nodes are also part of the node features. Another common feature type is related to the textual content represented by the node, e.g., word embeddings. The approaches differ in terms of which embeddings models are used. In this regard, BERT-based-embeddings are often employed, which was also the observation regarding textual encoders of sequence-based approaches. Some other methods are Byte Pair Encoding (Sennrich et al., 2016), BiLSTM and LayoutLM. In some cases, multiple text embeddings are used simultaneously. For example, (Zhang et al., 2021) employ Word2Vec, BERT and LayoutLM for the word embeddings. On top of positional and textual embeddings, many approaches also integrate visual features, often obtained from models such as ResNet or LayoutLM. Many authors also consider hand-crafted miscellaneous features. The goal of these features is usually to incorporate certain node characteristics as well as relationships between nodes that should aid the model in performing KIE. (Lohani et al., 2019) for example include various boolean features whether a graph node represents a date, zipcode or known city, among others. The approaches also differ regarding the size of the node feature vectors. Although in many cases the actual number is not explicitly reported in the manuscripts, on average, the features vectors have a length of around 500. In general, there is little discussion about embeddings sizes for node features in related work, from which it can be concluded that these have no significant influence on the performance of the KIE systems. Much more important seem to be the different types of features which are being integrated. Another distinguishing factor of graph-based approaches is, whether they incorporate a global node in the defined graph representation. This global node can contain information about the entire document image and can be connected to the individual nodes of the graph. To this end, only the approaches by (Hua et al., 2020; Zhang et al., 2023c) use such global nodes. (Liao et al., 2023) also construct a global node, however it is not used as a information carrier and is rather formally required due to the chosen concept based on document layout trees.
@@
Edges represent the second integral part of the graph structures. It is noticeable that a large portion of work does not incorporate any edge features at all. Consequently, less attention is paid to this type of feature compared to node embeddings. However, (Biescas et al., 2024) show that edge features, especially in terms of geometric data, can improve KIE performance, even more so than geometric information obtained from node features, as it integrates spatial information of the surroundings. The edge-features are usually derived from pairwise positional relationships between the connected nodes. Most often, horizontal and/or vertical distances, both in absolute and relative values, are considered. In this regard, (Liu et al., 2019a) highlight the importance of the visual distances between two nodes. The aspect ratio of the bounding boxes that span the graph nodes is also often included as edge features. In only three of the graph-based approaches are visual embeddings part of the edge features. Regarding the distribution of edge directions, i.e., whether the edges are directed or undirected, there is no clear dominant variant, however in the majority of the cases, the edges are undirected. Additional research is required to appropriately estimate the impact of certain design features such as the edge direction on the extraction performance. To this end, (Gbada et al., 2024a) show that both directed edges as well as introducing edge weights can improve extraction results. However, as illustrated before, these considerations are usually dependent on the underlying scenario and generalizing statements can only be made to a limited extent.